OpenStudy (anonymous):

Need help! Discriminant of 3x^2-2x+9=0?

OpenStudy (anonymous):

Discriminant... So 0, correct?

OpenStudy (anonymous):

\[b^2-4ac\] \[a=3,b=-2,c=9\] compute

OpenStudy (experimentx):

b^-4ac (-2)^2-4x3x9

OpenStudy (ash2326):

We have \[3x^2-2x+9=0\] Standard quadratic equation \[ax^2+bx+c=0 \] \[Discriminant(D)=b^2-4ac\] Here b=-2 a=3 c=9 so \[D=(-2)^2-4 \times 3 \times 9=4-4 \times 3 \times 9=-104\]

OpenStudy (anonymous):

(104)^.5i (imaginary).

OpenStudy (anonymous):

Whoa, thanks guys. You're awesome